# MFCPL: Multimodal Federated Cross Prototype Learning for Missing Modalities and Data Heterogeneity

### v1

@@ Abstract
Multimodal federated learning (MFL) has emerged as a decentralized machine learning paradigm, allowing multiple clients with different modalities to collaborate on training a machine learning model across diverse data sources without sharing their private data. However, challenges, such as data heterogeneity and severely missing modalities, pose crucial hindrances to the robustness of MFL, significantly impacting the performance of global model. The absence of a modality introduces misalignment during the local training phase, stemming from zero-filling in the case of clients with missing modalities. Consequently, achieving robust generalization in global model becomes imperative, especially when dealing with clients that have incomplete data. In this paper, we propose Multimodal Federated Cross Prototype Learning (MFCPL), a novel approach for MFL under severely missing modalities by conducting the complete prototypes to provide diverse modality knowledge in modality-shared level with the cross-modal regularization and modality-specific level with cross-modal contrastive mechanism. Additionally, our approach introduces the cross-modal alignment to provide regularization for modality-specific features, thereby enhancing overall performance, particularly in scenarios involving severely missing modalities. Through extensive experiments on three multimodal datasets, we demonstrate the effectiveness of MFCPL in mitigating these challenges and improving the overall performance.
@@ 1 Introduction
Figure 1: Problem Illustration of Multimodal Federated Learning. We present two challenges of MFL which are a) Missing Modality and b) Data Heterogeneity.
@@
Federated learning (FL) McMahan et al. (2017); Kairouz et al. (2021); Tan et al. (2022a) has emerged as a promising distributed machine learning paradigm, allowing multiple clients to collaboratively train models without leaking private data. Despite promising outcomes observed across various domains Rieke et al. (2020); Nguyen et al. (2021), most FL methods are tailored for scenarios with clients processing unimodal data, potentially constraining their applicability in real-world settings. However, in the era of data proliferation, users have the distinctive opportunity to harness many data types, unlocking new dimensions to explore the potential of multimodal data that aligns more closely with real-world applications Caesar et al. (2020); Fu et al. (2023). Therefore, with the necessity of handling multimodal data while preserving privacy, multimodal federated learning (MFL) has emerged as a promising paradigm Xiong et al. (2022); Zhao et al. (2022). In traditional FL scenarios, data heterogeneity Li et al. (2020); Deng et al. (2020), where distributed data is non-IID (identically and independently distributed), has become a crucial challenge, leading to performance degradation. Despite numerous studies in FL domain Luo et al. (2021); Mendieta et al. (2022) aimed at mitigating data heterogeneity, their predominant focus has been on addressing this challenge in applications involving unimodal data, rather than extending these solutions to the context of MFL. Additionally, in practical applications, it is essential to recognize that real-world scenarios may not always allow the utilization of all available modalities. Challenges such as sensor disconnections, different hardware platforms, or other operational issues can impede the efficient integration of diverse data streams, impacting the model’s robustness. While data heterogeneity presents its own set of challenges, the situation is compounded by the significant impact of severely missing modalities, leading to a further degradation in the model’s performance within the context of MFL Chen and Zhang (2022); Feng et al. (2023). Consequently, we pose a research question: How can we effectively address both data heterogeneity and missing modality challenges in MFL?. In Figure 1, we illustrate the concepts of data heterogeneity and missing modalities.
@@
In the context of MFL, when confronted with significant missing modalities, a common practice is to impute the missing data with zeros. However, this zero-filling strategy introduces potential misalignments or undesired effects, adversely impacting the model performance in multimodal learning. To enhance representation learning in MFL under conditions of missing modalities, an effective guiding method during training is crucial. Recently, prototype learning has been introduced as an effective information carrier between the server and clients in FL Tan et al. (2022c); Qiao et al. (2023). A prototype is defined as the mean value of features from the same class and is represented as a vector. Several studies in FL have leveraged prototypes to tackle issues such as data heterogeneity Tan et al. (2022b) or domain shift Huang et al. (2023). Nonetheless, prior research on FL with prototypes has solely focused on scenarios involving unimodal data. Building upon the inspiration derived from prototype learning, we aim to extend its application to address the challenge of missing modalities in MFL. To tackle the challenges of both data heterogeneity and the existence of missing modalities, we derive the complete prototype by aggregating local prototypes from various clients, including those with both missing and full modalities, utilizing the shared projection head. This ensures that the features of each class encapsulate knowledge from diverse modalities. Additionally, the complete prototype serves as guidance for clients experiencing missing modalities.
@@
In this paper, we propose Multimodal Federated Cross Prototype Learning (MFCPL), which consists of three main components: Cross-Modal Prototypes Regularization (CMPR), Cross-Modal Prototypes Contrastive (CMPC) and Cross-Modal Alignment (CMA). Drawing inspiration from the success of prototypes in addressing data heterogeneity in FL Tan et al. (2022b); Mu et al. (2023), we introduce a novel complete prototype to effectively handle missing modalities in MFL. First, we propose Cross-Modal Prototypes Regularization (CMPR) by shortening the distance between complete prototype and the local modality-shared representation, bridging the gap between local and global models. Through this process, diverse modality knowledge is effectively transferred from complete prototypes to those clients with missing modalities, thereby contributing to the robustness of the learning system under conditions of missing modalities. Second, we introduce Cross-Modal Prototypes Contrastive (CMPC), leveraging complete prototypes to facilitate cross-modal contrastive learning. CMPC encourages local modality-specific representations to be close to the complete prototypes within the same semantic class and pushes them away from complete prototypes of different classes. Cross-modal contrastive learning involves aligning modality-specific representations and complete prototypes into the same latent space through a projection layer. We can provide robust and rich semantic representations suitable for clients with missing modalities by aligning the complete prototype with modality-specific representations. Finally, we introduce Cross-Modal Alignment (CMA), which aligns the embedding features of the missing modality with those of the existing modality through a projection layer. By incorporating this, we mitigate the negative impact of missing modality and improve the generalization performance of the global model. These components collectively empower our proposed framework as a robust method to tackle severe missing modalities in heterogeneous MFL effectively. Our primary contributions are:
@@ 2.1 Multimodal Federated Learning
Numerous research works in multimodal federated learning (MFL) Xiong et al. (2022); Zhao et al. (2022); Yu et al. (2023), have explored diverse techniques to facilitate the decentralized training of machine learning models across multiple data sources. For instance, Xiong et al. Xiong et al. (2022) introduced a co-attention layer to merge representations from various modalities, obtaining global features to train personalized models for clients. Addressing labeled data constraints in MFL, Zhao et al. Zhao et al. (2022) proposed a semi-supervised framework within the FedAvg mechanism McMahan et al. (2017). CreamFL Yu et al. (2023) introduced a knowledge distillation-based MFL framework using a contrastive method for representation aggregation leveraging the public dataset. However, these methods do not consider the missing modality scenario in the system, limiting their practical applicability in real-world settings. To address the issue of missing modalities in MFL, Chen and Zhang Chen and Zhang (2022) employed a dynamic graph structure to capture correlations among different types of clients. Recently, FedMultimodal Feng et al. (2023) introduced the first benchmark in MFL that evaluates MFL under different types of data corruption: missing modalities, missing labels, and erroneous labels. In this paper, to address the missing modality problem in heterogeneous MFL, we introduce the novel concept of leveraging a complete prototype to operate with local representations at two levels: modality-specific representations and modality-shared representation through a projection layer.
@@ 2.2 Incomplete Multimodal Learning
Exploring incomplete data, such as missing modalities in multimodal learning, is crucial while ensuring robustness in real-world scenarios with missing modalities. Instances of missing modalities are prevalent in practical situations, often stemming from hardware constraints or storage limitations. Various approaches, such as data imputation Tran et al. (2017) or data reconstruction Pham et al. (2019); Zhao et al. (2021), have been proposed to address these challenges by filling in missing modalities based on existing ones. For instance,  Tran et al. Tran et al. (2017) proposed a cascaded residual autoencoder (CRA) to impute missing modalities. Additionally, Dong et al. Dong et al. (2023) introduced a cross-modal translation method to replace null entries in the embedding for missing modality generalization during the inference stage. Another approach involves utilizing a knowledge transfer mechanism to convey information from modalities with available data to those with missing data Garcia et al. (2018); Luo et al. (2018)Poklukar et al. Poklukar et al. (2022) introduced the Geometric Multimodal Contrastive (GMC) framework, which geometrically aligns complete and modality-specific representations to address missing modality information across various tasks. In this study, we employ the definition of a complete prototype to provide diverse modality knowledge for clients with missing modalities, thereby enhancing the local training process for these clients.
@@ 2.3 Prototype Learning
The prototype is computed as the mean feature vectors within the same class Snell et al. (2017) and has found applications in various research domains, including few-shot learning Xu et al. (2020); Zhang et al. (2021), and unsupervised learning Ye et al. (2019); Guo et al. (2022). In FL, prototypes have been employed in multiple studies to address data heterogeneity Tan et al. (2022b, c); Huang et al. (2023); Qiao et al. (2023). FedProto Tan et al. (2022b) was among the pioneering works to apply prototypes in FL, introducing a communication approach between clients and the server using prototypes instead of parameter exchange to learn personalized models. Additionally,  Qiao et al. Qiao et al. (2023) proposed the multi-prototype approach with the contrastive learning strategy to address the feature heterogeneity. Recently, FPL Huang et al. (2023) proposed the cluster prototypes to tackle the domain shift in FL. However, the majority of prior works concentrate on scenarios where clients exclusively possess unimodal data. In the context of MFL under missing modalities, it is crucial to address the generalization across various types of clients with diverse modality absences. Our work illuminates the use of the complete prototype to accomplish this objective at two levels: modality-specific and modality-shared representations.
@@ 3.1 Preliminaries
Multimodal federated learning (MFL) aims to facilitate collaborative training among multiple clients on a shared global model, leveraging multimodal data while ensuring the data privacy. In this paper, we consider that there are $N$ clients (indexed by $i$) with private multimodal data $D_{i}=\{{x}^{A}_{i},{x}^{B}_{i},{y}_{i}\}$, where $x^{A}_{i}$ and $x^{B}_{i}$ represent samples from modalities A and B, respectively, and $y_{i}$ denotes the label associated with each sample.
Figure 2: Illustration of MFCPL. Different client types upload their prototypes to server using projection head $g_{1}$. We provide diverse modality knowledge in modality-shared level with $\mathcal{L}_{CMPR}$ and modality-specific level with $\mathcal{L}_{CMPC}$ using projection head $g_{2}$. The alignment loss $\mathcal{L}_{CMA}$ is introduced to reduce the noise in with between projected modality-specific representations of two modalities using projection head $g_{2}$.
Client models share the identical architecture, comprising three modules: encoders $f_{m}$, self-attention fusion module $h_{\text{self-attention}}$, and classifier $g$. The encoders take input samples $x^{A}$ and $x^{B}$ and encode them into $n$-dimensional feature vectors $z_{A}=f_{A}(x^{A}),z_{B}=f_{B}(x^{B})\in\mathbb{R}^{n}$. Subsequently, a self-attention fusion module $h_{\text{self-attention}}$ integrates the representations from two modalities into the same feature space, denoted as $e=h_{\text{self-attention}}(z_{A},z_{B})\in\mathbb{R}^{u}$. Finally, the classifier $g$ maps the multimodal feature $e$ to the logits output $z_{cls}=g(e)\in\mathbb{R}^{I}$. Specifically, given the model parameters for the whole network as $\theta$, and $D$ is the sum of samples over all clients, we formulate the global objective:
$\operatorname*{argmin}_{\theta}{L}(\theta)$ $=\sum^{N}_{i=1}\frac{|D_{i}|}{|D|}{\mathcal{L}}_{i}(\theta_{i},D_{i}).$ (1)
where, the loss function $\mathcal{L}_{i}$ is the task training loss $\mathcal{L_{S}}=L(z_{cls},y)$ for $i^{th}$ client.

### v2

# Cross-Modal Prototype based Multimodal Federated Learning under Severely Missing Modality

@@ Abstract
Multimodal federated learning (MFL) has emerged as a decentralized machine learning paradigm, allowing multiple clients with different modalities to collaborate on training a global model across diverse data sources without sharing their private data. However, challenges, such as data heterogeneity and severely missing modalities, pose crucial hindrances to the robustness of MFL, significantly impacting the performance of global model. The occurrence of missing modalities in real-world applications, such as autonomous driving, often arises from factors like sensor failures, leading knowledge gaps during the training process. Specifically, the absence of a modality introduces misalignment during the local training phase, stemming from zero-filling in the case of clients with missing modalities. Consequently, achieving robust generalization in global model becomes imperative, especially when dealing with clients that have incomplete data. In this paper, we propose Multimodal Federated Cross Prototype Learning (MFCPL), a novel approach for MFL under severely missing modalities. Our MFCPL leverages the complete prototypes to provide diverse modality knowledge in modality-shared level with the cross-modal regularization and modality-specific level with cross-modal contrastive mechanism. Additionally, our approach introduces the cross-modal alignment to provide regularization for modality-specific features, thereby enhancing the overall performance, particularly in scenarios involving severely missing modalities. Through extensive experiments on three multimodal datasets, we demonstrate the effectiveness of MFCPL in mitigating the challenges of data heterogeneity and severely missing modalities while improving the overall performance and robustness of MFL.
keywords:
Federated learning, Multimodal learning; Representation learning; Prototype learning
[…]
@@ 1 Introduction
Federated learning (FL) (McMahan et al., 2017; Kairouz et al., 2021; Tan et al., 2022a) has emerged as a promising distributed machine learning paradigm, allowing multiple clients to collaboratively train models without leaking private data. Despite promising outcomes observed across various domains (Rieke et al., 2020; Nguyen et al., 2021), most FL methods are tailored for scenarios with clients processing unimodal data, potentially constraining their applicability in real-world settings. However, in the era of data proliferation, users have the distinctive opportunity to harness many data types, unlocking new dimensions to explore the potential of multimodal data that aligns more closely with real-world applications (Caesar et al., 2020; Fu et al., 2023). Therefore, with the necessity of handling multimodal data while preserving privacy, multimodal federated learning (MFL) has emerged as a promising paradigm (Xiong et al., 2022; Zhao et al., 2022). In traditional FL scenarios, data heterogeneity (Li et al., 2020; Deng et al., 2020), where distributed data is non-IID (identically and independently distributed), has become a crucial challenge, leading to performance degradation. Despite numerous studies in FL domain (Luo et al., 2021; Mendieta et al., 2022) aimed at mitigating data heterogeneity, their predominant focus has been on addressing this challenge in applications involving unimodal data, rather than extending these solutions to the context of MFL. Additionally, in practical applications, it is essential to recognize that real-world scenarios may not always allow the utilization of all available modalities, particularly in MFL, where the access to complete modalities by clients is often not guaranteed. For example, if activities are recorded using two sensors and one experiences a disconnection (e.g., Bluetooth connectivity issues in human activity recognition or malfunctioning camera sensors in autonomous driving), it leads to a missing modality. Considering another scenario in text and image classification, some clients may only have access to image data (e.g., image without the caption), while others may possess both image and accompanying image data (e.g., images uploaded along with a caption). Such missing modalities could lead to a significant performance degradation, as the model trained under incomplete modality condition fails to capture the knowledge of multimodal features. For example, on different multimodal datasets, FedAvg experiences a significant decrease in performance when the missing rate increases from $q=0.5$ to $q=1.0$, as shown in Figure 2. Challenges such as sensor disconnections, different hardware platforms, or other operational issues can impede the efficient integration of diverse data streams, impacting the model’s robustness.
Figure 1: Problem Illustration of Multimodal Federated Learning with $M=2$. We present two challenges of MFL which are a) Missing Modality and b) Data Heterogeneity.
While data heterogeneity presents its own set of challenges, the situation is compounded by the significant impact of severely missing modalities, leading to a further degradation in the model’s performance within the context of MFL (Chen and Zhang, 2022; Feng et al., 2023). Consequently, we pose a research question: How can we effectively address both data heterogeneity and missing modality challenges in MFL?. In Figure 1, we illustrate the concepts of data heterogeneity and missing modalities. Following FedMultimodal (Feng et al., 2023), we set the equal missing rate for each modality on each client, creating a scenario of missing modalities. We also experiment with scenarios where clients have partial access to data from missing modalities. Instead of assuming that missing modalities are completely unavailable, we consider cases where clients [^1] Individual entities (e.g, devices or nodes.) in federated learning system that own local data can access a specified proportion of the missing data, such as 20% or 50%. For example, in text and image classification task, clients may gain access to a designated portion of missing text data or image data. Similarly, when multiple accelerometer sensors are deployed, and some are disconnected in human activity recognition, clients can still utilize the data from the operational sensors. We will further this experimental setting in subsection 4.2.
Figure 2: Effect of different missing modality rates $q$ on FedAvg on different datasets.
@@
When dealing with significant missing modalities, a common practice is to impute the missing data with zeros. However, this zero-filling strategy introduces potential misalignments or undesired effects, adversely impacting the model performance in multimodal learning. To enhance representation learning in MFL under conditions of missing modalities, an effective guiding method during training is crucial. Recently, prototype learning has been introduced as an effective information carrier between the server and clients in FL (Tan et al., 2022c; Qiao et al., 2023). A prototype is defined as the mean value of features from the same class, represented as a vector. Prototypes are constructed by grouping feature vectors based on class labels and calculating their mean within a shared embedding space. This results in a compact and representative feature for each class in a low-dimensional feature space. Prototypes naturally reduce noise and outliers by averaging the feature vectors within a class, leading to robust and stable representations. Furthermore, they serve as lightweight class representatives by summarizing the meaningful core features of a class into a single compact vector, significantly reducing data size while retaining the most salient information  (Snell et al., 2017; Tan et al., 2022b; Zhang et al., 2024). Several studies in FL have leveraged prototypes to tackle issues such as data heterogeneity (Tan et al., 2022b) or domain shift (Huang et al., 2023). Nonetheless, prior research on FL with prototypes has solely focused on scenarios involving unimodal data. Building upon the inspiration derived from prototype learning, we aim to extend its application to address the challenge of missing modalities in MFL. To tackle the challenges of both data heterogeneity and the existence of missing modalities, we derive the complete prototype by aggregating local prototypes from various clients, including those with both missing and full modalities, utilizing the shared projection head. This ensures that the features of each class encapsulate knowledge from diverse modalities. Additionally, the complete prototype serves as guidance for clients experiencing missing modalities.
@@
In this paper, we propose Multimodal Federated Cross Prototype Learning (MFCPL), which consists of three main components: Cross-Modal Prototypes Regularization (CMPR), Cross-Modal Prototypes Contrastive (CMPC) and Cross-Modal Alignment (CMA). Drawing inspiration from the success of prototypes in addressing data heterogeneity in FL (Tan et al., 2022b; Mu et al., 2023), we introduce a novel complete prototype to effectively handle missing modalities in MFL. First, we propose Cross-Modal Prototypes Regularization (CMPR) by shortening the distance between complete prototype and the local modality-shared representation, bridging the gap between local and global models. Through this process, diverse modality knowledge is effectively transferred from complete prototypes to those clients with missing modalities, thereby contributing to the robustness of the learning system under conditions of missing modalities. Second, we introduce Cross-Modal Prototypes Contrastive (CMPC), leveraging complete prototypes to facilitate cross-modal contrastive learning. CMPC encourages local modality-specific representations to be close to the complete prototypes within the same semantic class and pushes them away from complete prototypes of different classes. Cross-modal contrastive learning involves aligning modality-specific representations and complete prototypes into the same latent space through a projection layer. We can provide robust and rich semantic representations suitable for clients with missing modalities by aligning the complete prototype with modality-specific representations. Finally, we introduce Cross-Modal Alignment (CMA), which aligns the embedding features of the missing modality with those of the existing modality through a projection layer. By incorporating this, we mitigate the negative impact of missing modality and improve the generalization performance of the global model. These components collectively empower our proposed framework as a robust method to tackle severe missing modalities in heterogeneous MFL effectively. The details of each component are provided in subsection 3.3. Our primary contributions are:
  • 1.
@@
Figure 3: Illustration of MFCPL with $M=2$. Different client types upload their local prototypes based on Eq. 3 to server using projection head $g_{1}$. We provide diverse modality knowledge with the complete prototypes from Eq. 4 in modality-shared level with $\mathcal{L}_{CMPR}$ from Eq. 5 and modality-specific level with $\mathcal{L}_{CMPC}$ from Eq. 8 using projection head $g_{2}$. The alignment loss $\mathcal{L}_{CMA}$ in Eq. 9 is introduced to reduce the negative effect arising from zero-filling of missing modalities and enhance the coherence between projected modality-specific representations of two modalities using projection head $g_{2}$. In the illustration, the square class represents the positive class, while the triangle class represents the negative class.
@@ 2.1 Multimodal Federated Learning
Numerous research works in multimodal federated learning (MFL) (Xiong et al., 2022; Zhao et al., 2022; Yu et al., 2023), have explored diverse techniques to facilitate the decentralized training of machine learning models across multiple data sources. For instance, (Xiong et al., 2022) introduced a co-attention layer to merge representations from various modalities, obtaining global features to train personalized models for clients. In a different approach, the authors in (Yang et al., 2022, 2024) addressed the task of cross-modal federated human activity recognition (HAR) by proposing a feature disentangled activity recognition network to capture modality-discriminative characteristics for each client. However, their work focuses on a cross-modal setting with HAR task, where each client possesses data from only one modality. In contrast, multimodal federated learning considers scenarios where clients may have access to multiple modalities. In our work, by adjusting the missing rate, we extend our approach to accommodate both cross-modal and multimodal setting. Addressing labeled data constraints in MFL, (Zhao et al., 2022) proposed a semi-supervised framework within the FedAvg mechanism (McMahan et al., 2017). CreamFL (Yu et al., 2023), FedMEKT (Le et al., 2024) introduced a knowledge distillation-based MFL framework for representation aggregation leveraging the public dataset. However, these above method rely on additional public datasets, which is impractical in real-world scenarios. Moreover, these methods did not consider the missing modality scenario in the system, limiting their practical applicability in real-world settings. To address the issue of missing modalities in MFL, (Chen and Zhang, 2022) employed a dynamic graph structure to capture correlations among different types of clients. Recently, FedMultimodal (Feng et al., 2023) introduced the first benchmark in MFL that evaluates MFL under different types of data corruption: missing modalities, missing labels, and erroneous labels. In this paper, to address the missing modality problem in heterogeneous MFL, we introduce the novel concept of leveraging a complete prototype to operate with local representations at two levels: modality-specific representations and modality-shared representation through a projection layer.
@@ 2.2 Incomplete Multimodal Learning
Exploring incomplete data, such as missing modalities in multimodal learning, is crucial while ensuring robustness in real-world scenarios with missing modalities. Instances of missing modalities are prevalent in practical situations, often stemming from hardware constraints or storage limitations. SMIL (Ma et al., 2021) proposed a Bayesian meta-learning framework that incorporates the missing modality reconstruction and feature regularization to tackle the problem of multimodal learning with severely missing modalities. Moreover, various approaches, such as data imputation (Tran et al., 2017) or data reconstruction (Pham et al., 2019; Zhao et al., 2021), have been proposed to address these challenges by filling in missing modalities based on existing ones. For instance, (Tran et al., 2017) proposed a cascaded residual autoencoder (CRA), which utilized stacked residual autoencoders to learn the complex relationships between different modalities to impute missing ones. Regarding model robustness, the authors in (Ma et al., 2022) investigate the Transformer robustness against modal-incomplete data and improve it via multi-task optimization. Additionally, (Dong et al., 2023) introduced a cross-modal translation method to replace null entries in the embedding for missing modality generalization during the inference stage. Another approach involves utilizing a knowledge transfer mechanism to convey information from modalities with available data to those with missing data (Garcia et al., 2018; Luo et al., 2018)(Poklukar et al., 2022) introduced the Geometric Multimodal Contrastive (GMC) framework, which geometrically aligns complete and modality-specific representations to address missing modality information across various tasks. In this study, we employ the definition of a complete prototype to provide diverse modality knowledge for clients with missing modalities, thereby enhancing the local training process for these clients.
@@ 2.3 Prototype Learning
The prototype is computed as the mean feature vectors within the same class (Snell et al., 2017) and has found applications in various research domains, including few-shot learning (Xu et al., 2020; Zhang et al., 2021), and unsupervised learning (Ye et al., 2019; Guo et al., 2022). Prototypes provide robust class representations by reducing noise and mitigating the influence of outliers. In FL, prototypes have been employed in multiple studies to address data heterogeneity (Tan et al., 2022b, c; Huang et al., 2023; Qiao et al., 2023). FedProto (Tan et al., 2022b) was among the pioneering works to apply prototypes in FL, introducing a communication approach between clients and the server using prototypes instead of parameter exchange to learn personalized models. Additionally,  (Qiao et al., 2023) proposed the multi-prototype approach with the contrastive learning strategy to address the feature heterogeneity. Recently, FPL (Huang et al., 2023) proposed the cluster prototypes to tackle the domain shift in FL. However, the majority of prior works concentrate on scenarios where clients exclusively possess unimodal data. In the context of MFL under missing modalities, it is crucial to address the generalization across various types of clients with diverse modality absences. Our work illuminates the use of the complete prototype to accomplish this objective at two levels: modality-specific and modality-shared representations.
@@ 3.1 Preliminaries
Multimodal federated learning (MFL) aims to facilitate collaborative training among multiple clients on a shared global model, leveraging multimodal data while ensuring data privacy. In this paper, we consider that there are $N$ clients (indexed by $i$) with private multimodal data $D_{i}=\{{x}^{1}_{i},{x}^{2}_{i},...,x^{M}_{i},{y}_{i}\}$, where $x^{m}_{i}$ represents samples from the modality $m\in M$ and $y_{i}$ denotes the label associated with each sample. In scenarios with missing modalities, the combination of $M$ modalities leads to $12^{M}-$ 1 possible types of clients, as each client may possess data corresponding to a unique subset of modalities. For example, for applications with image and text modalities, there are three types of client: clients missing image data, clients missing text, and clients with both image and text data.
Client models share the identical architecture, comprising three modules: specific encoder of each modality $f_{m}$, self-attention fusion module $h_{\text{self-attention}}$, and classifier $g$. Each encoder $f_{m}$ processes input samples from modality $m$ to generate $n$-dimensional feature vectors $z_{m}=f_{m}(x^{m})\in\mathbb{R}^{n}$. Subsequently, these modality-specific representations are then integrated into a unified multimodal representation by the self-attention fusion module $h_{\text{self-attention}}$, denoted as $e=h_{\text{self-attention}}([z_{1},z_{2},...z_{M}])\in\mathbb{R}^{u}$. Here, $[z_{1},z_{2},...z_{M}]$ represents the concatenation of modality-specific representations. Finally, the classifier $g$ maps the multimodal feature $e$ to the logits output $z_{cls}=g(e)\in\mathbb{R}^{I}$. Specifically, given the model parameters for the whole network as $\theta$, and $D$ is the sum of samples over all clients, we formulate the global objective:
$\operatorname*{argmin}_{\theta}{L}(\theta)$ $=\sum^{N}_{i=1}\frac{|D_{i}|}{|D|}{\mathcal{L}}_{i}(\theta_{i},D_{i}),$ (1)
where the loss function $\mathcal{L}_{i}$ is the task training loss $\mathcal{L_{S}}=L(z_{cls},y)$ for $i^{th}$ client.